OpenStudy (anonymous):

how would i do this problem, Evaluate: 100(1 - 0.02)10.

OpenStudy (anonymous):

980

OpenStudy (nenadmatematika):

=(100-2)10=98*10=980

OpenStudy (anonymous):

100(0.98)(10) 100(9.8) 980

OpenStudy (anonymous):

\[100\times (.98)^{10}\]

OpenStudy (anonymous):

a calculator gives me 81.7 rounded

OpenStudy (anonymous):

yea i forgot to the ^10, but thank you guys!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

then disregard the 3 answers and refer to satellite's!

OpenStudy (anonymous):

first 3*